The length of a rectangular parking lot is 40 ft more than the width. if the parking lot is w feet wide, express its area as a function of the width.
Answer by checkley79(3341)

L=W+40
P=2L+2W
P=2(W+40)+2W
P=2W+80+2W
P=4W+80
4W=-80-P
W=-80/4-P/4
W=-20-P/4 ANS.